OpenStudy (anonymous):

Write the standard form of the line that passes through the given points. Include your work in your final answer. Given points: (3,1) and (-2,3)

OpenStudy (dan815):

find slope from the given 2 points by dividing change in y by the change in x

OpenStudy (dan815):

(x1,y1)=(3,1) (x2,y2)=(-2,3) y2-y1=3-1 x2-x1=-2-3 slope = (y2-y1)/(x2-x1) =(3-1)/(-2-3) = 2/-5 Slope=m =2/-5 and line with slope m and passing through a point (xo,yo) can be written in an equation in the form y-yo = m(x-xo) so a line with slope -2/5 and passing through the point (3,1) y-1=-2/5(x-3) you can simply this expression to write in terms of y=..... y=-2/5(x-3)+1 =-2/5 x - 2/5 * 3 + 1 =-2/5x -1/5 this is in y=mx+b form where m is the slope -2/5 and b is the y intercept which is -1/5